OpenStudy (anonymous):

Lim x->- infinity of [(-3x^3 +4x+1) / (2-x(x^2-1)]

OpenStudy (anonymous):

= \[\lim_{x \rightarrow \infty} \frac{-3x^3 +4x+1}{-x^3+x+2}\]

OpenStudy (anonymous):

Consider the dominate terms (terms with highest power. \[= \frac{-3}{-1}= 3\]